OpenStudy (anonymous):

5^2-2[(5-6)-(-1-6)]=?

OpenStudy (anonymous):

25 - 2[(-1) - (-7)] = ? 25 - 2[-1 + 7] = ? 25 -2[6] = ? 25 - 12 = 13